### v1

# MaskGCT: Zero-Shot Text-to-Speech with Masked Generative Codec Transformer

Abstract
Nowadays, large-scale text-to-speech (TTS) systems are primarily divided into two types: autoregressive and non-autoregressive. The autoregressive systems have certain deficiencies in robustness and cannot control speech duration. In contrast, non-autoregressive systems require explicit prediction of phone-level duration, which may compromise their naturalness. We introduce the Masked Generative Codec Transformer (MaskGCT), a fully non-autoregressive model for TTS that does not require precise alignment information between text and speech. MaskGCT is a two-stage model: in the first stage, the model uses text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and in the second stage, the model predicts acoustic tokens conditioned on these semantic tokens. MaskGCT follows the mask-and-predict learning paradigm. During training, MaskGCT learns to predict masked semantic or acoustic tokens based on given conditions and prompts. During inference, the model generates tokens of a specified length in a parallel manner. We scale MaskGCT to a large-scale multilingual dataset with 100K hours of in-the-wild speech. Our experiments demonstrate that MaskGCT achieves superior or competitive performance compared to state-of-the-art zero-shot TTS systems in terms of quality, similarity, and intelligibility while offering higher generation efficiency than diffusion-based or autoregressive TTS models. Audio samples are available at https://maskgct.github.io/.
1 Introduction
In recent years, large-scale zero-shot text-to-speech (TTS) systems (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024) have achieved significant improvements by scaling data and model size, including both autoregressive (AR) (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024; Shen et al., 2023; Ju et al., 2024; Le et al., 2024; Jiang et al., 2023) and non-autoregressive (NAR) models (Shen et al., 2023; Ju et al., 2024; Le et al., 2024; Jiang et al., 2023). However, these systems still exhibit some problems. Currently, AR-based TTS systems typically quantize speech into discrete tokens and then use language models to sequentially generate these tokens, which offer diverse prosody but also suffer from problems such as poor robustness and slow inference speed. NAR-based models, typically based on diffusion (Ju et al., 2024; Shen et al., 2023), flow matching (Le et al., 2024), or GAN (Jiang et al., 2023), require text and speech alignment information as well as the prediction of phone-level duration, resulting in a complex pipeline and producing more standardized but less natural speech.
Recently, masked generative transformers, a class of generative models, have achieved significant results in the fields of image (Chang et al., 2022, 2023; Li et al., 2023a), video Yu et al. (2023a, b), and audio (Garcia et al., 2023; Li et al., 2024; Ziv et al., 2024) generation, demonstrating potential comparable to or superior to autoregressive models or diffusion models. These models employ a mask-and-predict training paradigm and utilize iterative parallel decoding during inference. Some previous works have attempted to introduce masked generative models into the field of TTS. SoundStorm (Borsos et al., 2023) was the first attempt to use a masked generative transformer to predict multi-layer acoustic tokens extracted from SoundStream, conditioned on speech semantic tokens; however, it still needs to receive the semantic tokens of an AR model as input. Thus, SoundStorm is more of an acoustic model that converts semantic tokens into acoustic tokens and does not fully utilize the powerful generative potential of masked generative models. NaturalSpeech 3 (Ju et al., 2024) decomposes speech into discrete token sequences representing different attributes through special designs and generates tokens representing different attributes through masked generative models. However, NaturalSpeech 3 still requires speech-text alignment information to predict the phone-level duration.
In this work, we propose MaskGCT, a fully non-autoregressive model for text-to-speech synthesis that uses masked generative transformers without requiring text-speech alignment information. MaskGCT is a two-stage system; both stages are trained using the mask-and-predict learning paradigm. The first stage, the text-to-semantic (T2S) model, predicts masked speech semantic representation tokens within an in-context learning paradigm, using text token sequences and prompt speech semantic token sequences as the prefix, without requiring precise alignment information between text and speech. The second stage, the semantic-to-acoustic (S2A) model, utilizes semantic tokens to predict masked acoustic tokens extracted from an RVQ-based speech codec with prompt acoustic tokens. During inference, MaskGCT can generate speech of various specified lengths with high quality and naturalness, requiring only 25 iteration steps for using text to predict semantic tokens, which is significantly fewer than AR-based models. Another key aspect of our method is that we train a VQVAE (Van Den Oord et al., 2017) to quantify self-supervised speech learning features, rather than using k-means to obtain semantic tokens as in previous work. This approach minimizes the information loss of semantic features even with a single codebook. We also explore the scalability of our methods in tasks such as speech translation, voice conversion, emotion control, and speech content editing, demonstrating the potential of MaskGCT as a foundational model for speech generation.
Our experiments demonstrate that MaskGCT has achieved performance comparable to or superior to that of existing models in terms of speech quality, similarity, prosody, and intelligibility. Specifically, (1) It achieves comparable or better quality and naturalness than the ground truth speech across three benchmarks (LibriSpeech, SeedTTS test-en, and SeedTTS test-zh) in terms of CMOS. (2) It achieves human-level similarity between the generated speech and the prompt speech, with improvements of +0.017, -0.002, and +0.027 in SIM-O and +0.156, +0.319 and +0.244 in SMOS for LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively. (3) It achieves comparable intelligibility in terms of word error rate (WER) across the three benchmarks and demonstrates stability within a reasonable range of speech duration, which also indicates the diversity and controllability of the generated speech.
The contribution of our proposed MaskGCT can be summarized as follows: (1) We propose an innovative system for zero-shot TTS, which is based on non-autoregressive masked generative transformers. We also introduce a speech discrete semantic representation based on training a VQVAE on speech self-supervised representations. (2) MaskGCT achieves human-level similarity, naturalness, and intelligibility by scaling data to 100K hours of in-the-wild speech. MaskGCT also demonstrates the ability to control the total duration of generated speech, showcasing the controllability and diversity of speech produced by MaskGCT. (3) We investigate the scalability of MaskGCT across various tasks, including speech translation, voice conversion, emotion control, and speech content editing, utilizing zero-shot learning or post-training methods. This exploration underscores the potential of MaskGCT as a foundational model for speech generation.
2 Related Work
Large-scale TTS. Traditional TTS systems (Ren et al., 2020, 2019; Tan et al., 2024; Wang et al., 2017; Kim et al., 2021) are trained to generate speech from a single speaker or multiple speakers using hours of high-quality transcribed training data. Modern large-scale TTS systems (Kharitonov et al., 2023; Wang et al., 2023; Łajszczak et al., 2024; Kim et al., 2024; Peng et al., 2024; Anastassiou et al., 2024) aim to achieve zero-shot TTS (synthesizing speech for unseen speakers with speech prompts) by scaling both the model and data size. These systems can be mainly divided into AR-based and NAR-based categories. For AR-based systems: SpearTTS (Kharitonov et al., 2023) utilizes three AR models to predict semantic tokens from text, coarse-grained acoustic tokens from semantic tokens, and fine-grained acoustic tokens from coarse-grained tokens. VALLE (Wang et al., 2023) predicts the first layer of acoustic tokens extracted from EnCodec (Défossez et al., 2022) using an AR LLM, and the final layers with a NAR LLM. VoiceCraft (Peng et al., 2024) employs a single AR model to predict multi-layer acoustic tokens in a delayed pattern (Copet et al., 2024). BASETTS (Łajszczak et al., 2024) predicts novel speech codes extracted from WavLM features and uses a GAN model for waveform reconstruction. For NAR-based systems: NaturalSpeech 2 (Shen et al., 2023) employs latent diffusion to predict the latent representations from a codec model (Zeghidour et al., 2021). VoiceBox (Le et al., 2024) uses flow matching and in-context learning to predict mel-spectrograms, following a similar framework. MegaTTS (Jiang et al., 2023) utilizes a GAN to predict mel-spectrograms, while an AR model predicts phone-level prosody codes. NaturalSpeech 3 (Ju et al., 2024) employs a unified framework based on discrete diffusion models to predict discrete representations of different speech attributes. However, these NAR systems need to predict phoneme-level duration, leading to a complex pipeline and more standardized generative results. SimpleSpeech (Yang et al., 2024), DiTTo-TTS (Lee et al., 2024), and E2 TTS (Eskimez et al., 2024) are also NAR-based models that do not require precise alignment information between text and speech, nor do they predict phoneme-level duration. We discuss these concurrent works in Appendix L.
Masked Generative Model. Masked generative transformers, a class of generative models, achieve significant results and demonstrate potential comparable to or superior to that of autoregressive models or diffusion models in the fields of image (Chang et al., 2022, 2023; Lezama et al., 2022; Li et al., 2023a), video (Yu et al., 2023a, b), and audio (Borsos et al., 2023; Garcia et al., 2023; Li et al., 2024; Ziv et al., 2024) generation. MaskGIT (Chang et al., 2022) is the first work to use masked generative models for both unconditional and conditional image generation. Subsequently, Muse (Chang et al., 2023) leverages rich text to achieve high-quality and diverse text-to-image generation within the same framework. MAGVIT-v2 (Yu et al., 2023b) employs masked generative models with the novel lookup-free quantization (LFQ), outperforming diffusion models in image and video generation. Recently, some efforts have been made to adapt masked generative models to the field of audio generation. SoundStorm (Borsos et al., 2023) takes in the semantic tokens from AudioLM and utilizes this generative paradigm to generate tokens for a neural audio codec (Zeghidour et al., 2021). VampNet (Garcia et al., 2023) and MAGNeT (Ziv et al., 2024) apply masked generative models for music and audio generation, while MaskSR (Li et al., 2024) extends these models for speech restoration.
Discrete Speech Representation. Speech representation is a crucial aspect of speech generation. Early works (Ren et al., 2019; Wang et al., 2017) typically utilized mel-spectrograms as the modeling target. Recently, some large-scale TTS systems (Wang et al., 2023; Ju et al., 2024) have shifted to using discrete speech representations. Discrete speech representation can be primarily divided into two types: semantic discrete representation and acoustic discrete representation. Semantic discrete representations are mainly extracted from various speech SSL models (Chung et al., 2021; Hsu et al., 2021; Chen et al., 2022) using quantization methods such as k-means. Acoustic discrete representations, on the other hand, are usually obtained by training a VQGAN model (Van Den Oord et al., 2017) with the goal of waveform reconstruction, as seen in speech neural codecs (Défossez et al., 2022; Zeghidour et al., 2021; Kumar et al., 2024; Yang et al., 2023). Semantic discrete representation typically shows a stronger correlation with text, whereas acoustic discrete representation more effectively reconstructs audio. Consequently, some two-stage TTS models predict both semantic and acoustic tokens. FACodec (Ju et al., 2024) is a novel speech codec that disentangles speech waveforms into subspaces of different attributes, including content, prosody, timbre, and acoustic details.
3 Method
3.1 Background: Non-Autoregressive Masked Generative Transformer
Given a discrete representation sequence $\mathbf{X}$ of some data, we define $\mathbf{X}_{t}=\mathbf{X}\odot\mathbf{M}_{t}$ as the process of masking a subset of tokens in $\mathbf{X}$ with the corresponding binary mask $\mathbf{M}_{t}=[m_{t,i}]_{i=1}^{N}$. Specifically, this involves replacing $x_{i}$ with a special [MASK] token if $m_{t,i}=1$, and otherwise leaving $x_{i}$ unmasked if $m_{t,i}=0$. Here, each $m_{t,i}$ is independently and identically distributed according to a Bernoulli distribution with parameter $\gamma(t)$, where $\gamma(t)\in(0,1]$ represents a mask schedule function (for example, $\gamma(t)=\sin(\frac{\pi t}{2T}),t\in(0,T]$). We denote $\mathbf{X}_{0}=\mathbf{X}$. The non-autoregressive masked generative transformers are trained to predict the masked tokens based on the unmasked tokens and a condition $\mathbf{C}$. This prediction is modeled as $p_{\theta}(\mathbf{X}_{0}|\mathbf{X}_{t},\mathbf{C})$. The parameters $\theta$ are optimized to minimize the negative log-likelihood of the masked tokens:
$\mathcal{L}_{\text{mask}}$ $=\mathop{\mathbb{E}}\limits_{\mathbf{X}\in\mathcal{D},t\in\left[0,T\right]}-\sum_{i=1}^{N}m_{t,i}\cdot\log(p_{\theta}(x_{i}|\mathbf{X}_{t},\mathbf{C})).$
At the inference stage, we decode the tokens in parallel through iterative decoding. We start with a fully masked sequence $\mathbf{X}_{T}$. Assuming the total number of decoding steps is $S$, for each step $i$ from 1 to $S$, we first sample $\mathbf{\hat{X}}_{0}$ from $p_{\theta}(\mathbf{X}_{0}|\mathbf{X}_{T-(i-1)\cdot\frac{T}{S}},\mathbf{C})$. Then, we sample $\lfloor N\cdot\gamma(T-i\cdot\frac{T}{S})\rfloor$ tokens based on the confidence score to remask, resulting in $\mathbf{X}_{T-i\cdot\frac{T}{S}}$, where $N$ is the total number of tokens in $\mathbf{X}$. The confidence score for $\hat{x}_{i}$ in $\mathbf{\hat{X}}_{0}$ is assigned to $p_{\theta}(\mathbf{X}_{0}|\mathbf{X}_{T-(i-1)\cdot\frac{T}{S}},\mathbf{C})$ if $x_{T-(i-1)\cdot\frac{T}{S},i}$ is a [MASK] token; otherwise, we set the confidence score of $\hat{x}_{i}$ to $1$, indicating that tokens already unmasked in $\mathbf{X}_{T-(i-1)\cdot\frac{T}{S}}$ will not be remasked. Particularly, we choose $\lfloor N\cdot\gamma(T-i\cdot\frac{T}{S})\rfloor$ tokens with the lowest confidence scores to be masked.
The masked generative modeling paradigm was first introduced in Chang et al. (2022), and subsequent work such as Lezama et al. (2022) has further explored it under the perspective of discrete diffusion.
3.2 Model Overview
Following (Betker, 2023; Borsos et al., 2023; Wang et al., 2023), MaskGCT is a two-stage TTS system. The first stage uses text to predict speech semantic representation tokens, which contain most information of content and partial information of prosody. The second stage model is trained to learn more acoustic information. Unlike previous works (Betker, 2023; Wang et al., 2023; Borsos et al., 2023; Kharitonov et al., 2023) use an autoregressive model for the first stage, MaskGCT utilizes the non-autoregressive masked generative modeling paradigm for both the two stages without any alignment information between text and speech: (1) For the first stage model, we trained a model to learn $p_{\theta_{\text{s1}}}(\mathbf{S}|\mathbf{S}_{t},(\mathbf{S}^{p},\mathbf{P}))$, where $\mathbf{S}$ is the speech semantic representation token sequence obtained from a speech semantic representation codec (we introduce in 3.3), $\mathbf{S}^{p}$ is the prompt semantic token sequence, and $\mathbf{P}$ is the text token sequence. $\mathbf{S}^{p}$ and $\mathbf{P}$ are the condition for the first stage model. (2) The second stage model is trained to learn $p_{\theta_{\text{s2}}}(\mathbf{A}|\mathbf{A}_{t},(\mathbf{A}^{p},\mathbf{S}))$, where $\mathbf{A}$ is the multi-layer acoustic token sequence from a speech acoustic codec like (Zeghidour et al., 2021; Défossez et al., 2022). Our second stage model is similar to SoundStorm (Borsos et al., 2023). The overview of MaskGCT is presented in Figure 1. We give more details about the four parts in the following sections.
Figure 1: An overview of our MaskGCT system. MaskGCT consists of four main parts: (1) a speech semantic representation codec converts speech to semantic tokens; (2) a text-to-semantic model predicts semantic tokens with text and prompt semantic tokens; (3) a semantic-to-acoustic model predicts acoustic tokens conditioned on semantic tokens; (4) a speech acoustic codec reconstructs speech waveform from acoustic tokens.
3.3 Speech Semantic Representation Codec
Discrete speech representations can be divided into semantic tokens and acoustic tokens. Generally, semantic tokens are obtained by discretizing features from speech self-supervised learning (SSL). Previous two-stage, large-scale TTS systems (Betker, 2023; Borsos et al., 2023; Kharitonov et al., 2023) typically first use text to predict semantic tokens, and then employ another model to predict acoustic tokens or features. This is because semantic tokens have a stronger correlation with text or phonemes, which makes predicting them more straightforward than directly predicting acoustic tokens. Commonly, previous works have used k-means to discretize semantic features to obtain semantic tokens; however, this method can lead to a loss of information. This loss may complicate the accurate reconstruction of high-quality speech or the precise prediction of acoustic tokens, especially for tonally rich languages. For example, our early experiments demonstrate the challenges of accurately predicting acoustic tokens to achieve proper prosody for Chinese using semantic tokens obtained via k-means. Therefore, we need to discretize semantic representation features while minimizing information loss. Inspired by Huang et al. (2023), we train an encoder-decoder model to learn a vector quantization codebook that reconstructs speech semantic representations from a speech SSL model (e.g., W2v-BERT (Chung et al., 2021)). For a speech semantic representation sequence $\mathbf{S}\in\mathbb{R}^{T\times d}$, the vector quantizer quantizes the output of the encoder $\mathcal{E}(\mathbf{S})$ to $\mathbf{E}$, and the decoder reconstructs $\mathbf{E}$ back to $\hat{\mathbf{S}}$. We optimize the encoder and the decoder using a reconstruction loss between $\mathbf{S}$ and $\hat{\mathbf{S}}$, employ codebook loss to optimize the codebook and use commitment loss to optimize the encoder with the straight-through method (Van Den Oord et al., 2017). The total loss for training the semantic representation codec can be written as:
$\mathcal{L}_{\text{total}}$ $=\frac{1}{Td}(\lambda_{\text{rec}}\cdot||\mathbf{S}-\hat{\mathbf{S}}||_{1}+\lambda_{\text{codebook}}\cdot||\text{sg}(\mathcal{E}(\mathbf{S}))-\mathbf{E}||_{2}+\lambda_{\text{commit}}\cdot||\text{sg}(\mathbf{E})-\mathcal{E}(\mathbf{S})||_{2}).$
where sg means stop-gradient.
In detail, we utilize the hidden states from the 17th layer of W2v-BERT 2.0 (Chung et al., 2021) as the semantic features for our speech encoder. The encoder and decoder are composed of multiple ConvNext (Liu et al., 2022a) blocks. Following the methods of improved VQGAN (Yu et al., 2021) and DAC (Kumar et al., 2024), we use factorized codes to project the output of the encoder into a low-dimensional latent variable space. The codebook contains 8,192 entries, each of dimension 8. Further details about the model architecture are provided in the Appendix B.
3.4 Speech Acoustic Codec
Speech acoustic codec is trained to quantize speech waveform to multi-layer discrete tokens while aiming to preserve all the information of the speech as soon as possible. We follow the residual vector quantization (RVQ) method to compress the 24K sampling rate speech waveform into discrete tokens of 12 layers. The codebook size of each layer is 1,024 and the codebook dimension is 8. The model architectures, discriminators, and training losses follow DAC (Kumar et al., 2024), except that we use the Vocos (Siuzdak, 2023) architecture as the decoder for more efficient training and inference. Figure 5 shows the comparison between the semantic codec and acoustic codec.
3.5 Text-to-Semantic Model
Based on the previous discussion, we employ a non-autoregressive masked generative transformer to train a text-to-semantic (T2S) model, instead of using an autoregressive model or any text-to-speech alignment information. During training, we randomly extract a portion of the prefix of the semantic token sequence as the prompt, denoted as $\mathbf{S}^{p}$. We then concatenate the text token sequence $\mathbf{P}$ with $\mathbf{S}^{p}$ to form the condition. We simply add $(\mathbf{P},\mathbf{S}^{p})$ as the prefix sequence to the input masked semantic token sequence $\mathbf{S}_{t}$ to leverage the in-context learning ability of language models. We use a Llama-style (Touvron et al., 2023) transformer as the backbone of our model, incorporating gated linear units with GELU (Hendrycks and Gimpel, 2016) activation, rotation position encoding (Su et al., 2024), etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm (Zhang and Sennrich, 2019), which accepts the time step $t$ as the condition.
During inference, we generate the target semantic token sequence of any specified length conditioned on the text and the prompt semantic token sequence. In this paper, we also train a flow matching (Lipman et al., 2022) based duration prediction model to predict the total duration conditioned on the text and prompt speech duration, leveraging in-context learning. More details can be found in Appendix C.
3.6 Semantic-to-Acoustic Model
We also train a semantic-to-acoustic (S2A) model using a masked generative codec transformer conditioned on the semantic tokens. Our semantic-to-acoustic model is based on SoundStorm  (Borsos et al., 2023), which generates multi-layer acoustic token sequences. Given $N$ layers of the acoustic token sequence $\mathbf{A}^{1:N}$, during training, we select one layer $j$ from $1$ to $N$. We denote the $j$th layer of the acoustic token sequence as $A^{j}$. Following the previous discussion, we mask $A^{j}$ at the timestep $t$ to get $\mathbf{A}^{j}_{t}$. The model is then trained to predict $\mathbf{A}^{j}$ conditioned on the prompt $\mathbf{A}^{p}$, the corresponding semantic token sequence $\mathbf{S}$, and all the layers smaller than $j$ of the acoustic tokens. This can be formulated as $p_{\theta_{\text{s2a}}}(\mathbf{A}^{j}|\mathbf{A}^{j}_{t},(\mathbf{A}^{p},\mathbf{S},\mathbf{A}^{1:j-1}))$. We sample $j$ according to a linear schedule $p(j)=1-\frac{2j}{N(N+1)}$. For the input of the S2A model, since the number of frames in the semantic token sequence is equal to the sum of the frames in the prompt acoustic sequence and the target acoustic sequence, we simply sum the embeddings of the semantic tokens and the embeddings of the acoustic tokens from layer $1$ to $j$. During inference, we generate tokens for each layer from coarse to fine, using iterative parallel decoding within each layer. Figure 2 shows a simplified training diagram of the T2S and S2A models.
Figure 2: An overview of training diagram or the T2S (left) and S2A (right) models. The T2S model is trained to predict masked semantic tokens with text and prompt semantic tokens as the prefix. The S2A model is trained to predict masked acoustic tokens of a random layer conditioned on prompt acoustic tokens, semantic tokens, and acoustic tokens of the previous layers.
3.7 Other Applications
MaskGCT can accomplish tasks beyond zero-shot TTS, such as emotion control, speech content editing, voice conversion, and duration controllable speech translation with simple modifications or the assistance of external tools, demonstrating the potential of MaskGCT as a foundational model for speech generation. We provide more details in Appendix G,  H,  I,  J.
4 Experiments and Results
4.1 Experimental Settings
Datasets. We use the Emilia (He et al., 2024) dataset to train our models. Emilia is a multilingual and diverse in-the-wild speech dataset designed for large-scale speech generation. In this work, we use English and Chinese data from Emilia, each with 50K hours of speech (totaling 100K hours). We evaluate our zero-shot TTS models with three benchmarks: (1) LibriSpeech (Panayotov et al., 2015) test-clean, a widely used test set for English zero-shot TTS. (2) SeedTTS test-en, a test set introduced in Seed-TTS (Anastassiou et al., 2024) of samples extracted from English public corpora, specifically, the set includes 1,000 samples from the Common Voice dataset (Ardila et al., 2019). (3) SeedTTS test-zh, a test set introduced in Seed-TTS of samples extracted from Chinese public corpora, specifically, the set includes 2,000 samples from the DiDiSpeech dataset (Guo et al., 2021).
Evaluation Metrics. We use both objective and subjective metrics to evaluate our models. For the objective metrics, we evaluate speaker similarity (SIM), robustness (WER), and speech quality (FSD). Specifically, for speaker similarity, we compute the cosine similarity between the WavLM TDNN [^1] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) speaker embedding of generated samples and the prompt. For Word Error Rate (WER), we use a HuBERT-based [^2] https://huggingface.co/facebook/hubert-large-ls960-ft ASR model for LibriSpeech test-clean, Whisper-large-v3 for Seed-TTS test-en, and Paraformer-zh for Seed-TTS test-zh, following previous works. For speech quality, we use Fréchet Speech Distance (FSD) with self-supervised wav2vec 2.0 (Baevski et al., 2020) features, following Le et al. (2024). For the subjective metrics, comparative mean option score (CMOS) and similarity mean option score (SMOS) are used to evaluate naturalness and similarity, respectively. CMOS is on a scale of -3 to 3, and SMOS is on a scale of 1 to 5.
Baseline. We compare our models with state-of-the-art zero-shot TTS systems, including NaturalSpeech 3 (Ju et al., 2024), VALL-E (Wang et al., 2023), VoiceBox (Le et al., 2024), VoiceCraft (Peng et al., 2024), XTTS-v2 (Casanova et al., 2024), and CosyVoice (Du et al., 2024a). More details of each model can be found in Appendix F. We also train an AR T2S model to replace the T2S part of MaskGCT, which can be seen as the SoundStorm model for TTS.
Training. We train all models on 8 NVIDIA A100 80GB GPUs. We train two T2S models of different sizes (denoted as T2S-Base and T2S-large). For more details about the model architecture, please refer to Appendix A. We report the metrics of T2S-large by default, and you can find a comparison of model sizes in Section 4.3. We also compare two different methods of text tokenization: Grapheme-to-Phoneme(G2P) (Bernard and Titeux, 2021) and Byte Pair Encoding (BPE) (Gage, 1994). See more details of the two methods in Appendix D. We report the metrics of G2P by default. We optimize these models with the AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-4 and 32K warmup steps, following the inverse square root learning schedule. We use the classifier-free guidance (Ho and Salimans, 2022), during training for both the T2S and S2A models, we drop the prompt with a probability of 0.15. See more details about classifier-free guidance and classifier-free guidance rescale in Appendix E.
Inference. For the T2S model, we use 50 steps as the default total inference steps. The classifier-free guidance scale and the classifier-free guidance rescale factor (Lin et al., 2024) are set to 2.5 and 0.75, respectively. For sampling, we use a top-k of 20, with the sampling temperature annealing from 1.5 to 0. We add Gumbel noise to token confidences when determining the remasking process, following Chang et al. (2022). For the S2A model, we use $[40,16,1,1,1,1,1,1,1,1,1,1]$ steps for acoustic RVQ layers by default, we find the S2A model can also perform well with fewer inference steps of $[10,1,1,1,1,1,1,1,1,1,1,1]$. We use the same sampling strategy as the T2S model, except that we use greedy sampling instead of top-k sampling if the inference step is 1.
4.2 Zero-Shot TTS
Table 1: Evaluation results for MaskGCT and the baseline methods on LibriSpeech test-clean, SeedTTS test-en, SeedTTS test-zh. The boldface denotes the best result, the underline denotes the second best. gt length denotes the result obtained by using ground truth total speech length. The results in ‘()’ means the result is the best one selected from five random samples (rerank 5).
System SIM-O $\uparrow$ WER $\downarrow$ FSD $\downarrow$ SMOS $\uparrow$ CMOS $\uparrow$
LibriSpeech test-clean
Ground Truth 0.68 1.94 - 4.048 0.000
VALL-E (Wang et al., 2023) 0.50 5.90 - 3.472 -0.522
VoiceBox (Le et al., 2024) 0.64 2.03 0.762 3.802 -0.407
NaturalSpeech 3 (Ju et al., 2024) 0.67 1.94 0.786 4.258 0.164
VoiceCraft (Peng et al., 2024) 0.45 4.68 0.981 3.519 -0.333
XTTS-v2 (Casanova et al., 2024) 0.51 4.20 0.945 3.015 -0.985
MaskGCT 0.687(0.723) 2.634(1.976) 0.886 4.274 0.109
MaskGCT (gt length) 0.697 2.012 0.746 4.333 0.130
SeedTTS test-en
Ground Truth 0.730 2.143 - 3.923 0.000
CosyVoice (Du et al., 2024a) 0.643 4.079 0.316 3.516 -0.416
XTTS-v2 (Casanova et al., 2024) 0.463 3.248 0.484 3.150 -0.866
VoiceCraft (Peng et al., 2024) 0.470 7.556 0.226 3.179 -1.083
MaskGCT 0.717(0.760) 2.623(1.283) 0.188 4.242 0.033
MaskGCT (gt length) 0.728 2.466 0.159 4.129 0.122
SeedTTS test-zh
Ground Truth 0.750 1.254 - 3.863 0.000
CosyVoice (Du et al., 2024a) 0.750 4.089 0.276 3.544 -0.450
XTTS-v2 (Casanova et al., 2024) 0.635 2.876 0.413 2.950 -0.807
MaskGCT 0.774(0.805) 2.273(0.843) 0.106 4.094 0.050
MaskGCT (gt length) 0.777 2.183 0.101 4.107 0.083
In this section, we show the main results of zero-shot TTS: we show comparison results with SOTA baselines in Section 4.2.1; we compare MaskGCT with replacing T2S model to an AR model in Section 4.2.2; We present the performance of MaskGCT across varying speech tempos in Section 4.2.3.
4.2.1 Comparison with Baselines
We compare MaskGCT with baselines in terms of similarity, robustness, and generation quality. The main results are shown in Table 1. MaskGCT demonstrates excellent performance on all metrics and achieves human-level similarity, naturalness, and intelligibility. In similarity, MaskGCT’s SIM and SMOS both outperform the best baseline, whether assessed using the total length of ground truth or the predicted total duration (0.67$\rightarrow$0.687 in LibriSpeech, 0.643$\rightarrow$0.717 in SeedTTS test-en, 0.75$\rightarrow$0.774 in SeedTTS test-zh for SIM; +0.016 in LibriSpeech, +0.762 in SeedTTS test-en, +0.550 in SeedTTS test-zh for SMOS). When compared with human recordings, MaskGCT achieves human-level similarity across all three test sets (+0.017, -0.002, and +0.027 for SIM respectively in the three test sets, and +0.156, +0.319, and +0.244 for SMOS respectively in the three test sets). In robustness, MaskGCT likewise results nearly on par with ground truth (with 2.634, 2.623, 2.273 WER on LibriSpeech, SeedTTS test-en, and SeedTTS test-zh, respectively), exhibiting enhanced robustness compared to AR-based models and performing on par or better than NAR-based models such as VoiceBox and NaturalSpeech 3, without relying on phone-level duration predictions. In generation quality, MaskGCT achieves +0.109, +0.033, +0.050, and +0.083 CMOS across the three test sets when compared with human recordings, indicating that MaskGCT attains human-level naturalness on these test sets. We also observe that MaskGCT exhibits excellent performance when using both ground truth total duration and predicted total duration, indicating the robustness of MaskGCT within a reasonable range of total speech duration and the capability of our total duration predictor to yield appropriate durations.
4.2.2 Autoregressive vs. Masked Generative Models
We compare MaskGCT to replacing T2S MaskGCT with an AR T2S model (which we call AR + SoundStorm). Table 2 shows the performance of these two models on all three test sets. MaskGCT demonstrates improved similarity, robustness, and CMOS (+0.126 on LibriSpeech test-clean, +0.051 on SeedTTS test-en, and +0.323 on SeedTTS test-zh) across all three test sets. We also conducted comparisons on more challenging hard cases (such as repeating words, and tongue twisters, which are often considered as samples where TTS systems are prone to hallucinations). MaskGCT exhibited a more pronounced robustness advantage in these scenarios. See details in the Appendix K. In addition, compared to AR-based models, MaskGCT offers the capability to control the total duration of the generated speech, along with fewer inference steps, requiring only 25 to 50 steps for T2S models to achieve optimal results for speeches of any length. Conversely, the inference steps for AR-based models increase linearly with the length of the speech.
4.2.3 Duration Length Analysis
We analyze the robustness of the generated results of MaskGCT under different changes in total duration length (which can also be regarded as changes in speech tempo). The results are shown in Figure 3. We explore the results of multiplying the ground truth total duration by 0.7 to 1.3. The results show that the lowest WER is achieved at a total duration multiplier of 1.0, indicating that the models perform best when the speech is played at its natural speed. When the multiplier is 0.9 or 1.1, the model is still able to achieve a WER very close to the best. When the multiplier is 0.7 or 1.3, the WER is slightly higher but still within a reasonable range. This shows that our model can generate reasonable and accurate content at different speech tempos.
Table 2: Comparison results of the evaluation of MaskGCT and AR+SoundStorm. AR+SoundStorm can be regarded as replacing the T2S MaskGCT with the AR T2S model.
System SIM-O $\uparrow$ WER $\downarrow$ FSD $\downarrow$ SMOS $\uparrow$ CMOS $\uparrow$
LibriSpeech test-clean
AR + SoundStorm 0.672 3.267 0.998 4.203 -0.017
MaskGCT 0.687 2.634 0.886 4.274 0.109
SeedTTS test-en
AR + SoundStorm 0.683 2.846 0.323 4.026 -0.051
MaskGCT 0.717 2.623 0.188 4.242 0.000
SeedTTS test-zh
AR + SoundStorm 0.747 3.865 0.238 3.775 -0.323
MaskGCT 0.774 2.273 0.106 4.094 0.000
0.70.70.70.70.80.80.80.80.90.90.90.911111.11.11.11.11.21.21.21.21.31.31.31.3333344445555Total Duration MultiplierWERSeedTTS test-enSeedTTS test-zh
Figure 3: WER vs. Total Duration Multiplier (Speech Tempo) for SeedTTS test-en and SeedTTS test-zh test sets. We use inference timesteps as 25 for the T2S model.
4.3 Ablation Study
Inference Timesteps. Figure 4 shows the relationship between inference steps and metrics SIM and WER for SeedTTS test-zh (left) and test-en (right). Initially, SIM increases significantly, stabilizing after 25 steps. For test-zh, SIM rises from 0.761 at 5 steps to 0.771 at 75 steps, and for test-en, from 0.696 to 0.715. SIM reaches high values with just 10 steps but peaks around 25 steps. WER improves more dramatically, especially up to 25 steps. For test-zh, WER drops from 10.19 at 5 steps to 2.507 at 25 steps, and for test-en, from 8.096 to 2.346. Both SIM and WER show minimal changes beyond 25 steps. These findings indicate that while SIM metrics can be sufficiently optimized with around 10 inference steps, achieving the lowest WER values requires approximately 25 inference steps. Beyond this threshold, both SIM and WER metrics exhibit minimal changes, implying that further increases in inference steps do not yield substantial improvements in these performance metrics. Therefore, for practical applications, 25 inference steps may be considered optimal for balancing SIM and WER, ensuring efficient and effective performance.
5555101010102525252550505050757575750.750.750.750.750.760.760.760.760.770.770.770.770.780.780.780.780.790.790.790.790.80.80.80.8Inference StepsSIMSim22224444666688881010101012121212WERWER
5555101010102525252550505050757575750.680.680.680.680.70.70.70.70.720.720.720.720.740.740.740.74Inference StepsSIMSIM2222444466668888WERWER
Figure 4: Inference Steps vs. SIM and WER. The results on the left are for SeedTTS test-zh, and the results on the right are for SeedTTS test-en. In this ablation study, we utilize the ground truth speech length.
Model Size. We compare the performance differences of T2S models with varying model sizes. The result is shown in Table 3. We observe that the large model outperforms the base model across all metrics, albeit not significantly. We suggest that our system can achieve good performance with just the setting of the base model when using 100K hours of data. In the future, we will explore more comprehensive scaling laws for both model size and data scaling.
Table 3: Comparison results between T2S-Large and T2S-Base.
System SIM-O $\uparrow$ WER $\downarrow$ FSD $\downarrow$ #Parameters
SeedTTS test-en
T2S-Base 0.714 2.514 0.189 315M
T2S-Large 0.728 0.466 0.159 695M
SeedTTS test-zh
T2S-Base 0.769 2.216 0.123 315M
T2S-Large 0.777 2.183 0.101 695M
Text Tokenizer. We compare two different methods of text tokenization: (1) Grapheme-to-Phoneme(G2P), which converts text into a sequence of phonemes; (2) Byte Pair Encoding (BPE). See more details for the methods of text tokenization in Appendix D. The results indicate that G2P outperforms BPE in English with a higher similarity score (SIM) of 0.728 compared to 0.711 and a lower Word Error Rate (WER) of 2.466 versus 4.036. Conversely, in Chinese, G2P maintains a slightly higher similarity score (0.777 vs. 0.769) but BPE achieves a lower WER (1.921 vs. 2.338). These findings suggest that while G2P is superior in preserving text similarity and reducing errors in English, BPE is more effective in minimizing WER in Chinese. We hypothesize that the reason might be that the Chinese G2P system we used still has deficiencies in handling polyphonic characters, whereas BPE is capable of learning different pronunciations for the same character based on context.
Table 4: Text tokenizer: G2P vs. BPE. In this ablation study, we utilize the ground truth speech length.
SIM-O $\uparrow$ WER $\downarrow$
SeedTTS test-en
G2P 0.728 2.466
BPE 0.711 4.036
SeedTTS test-zh
G2P 0.777 2.183
BPE 0.769 1.921
5 Conclusion
In this paper, we present MaskGCT, a large-scale zero-shot TTS system that leverages fully non-autoregressive masked generative codec transformers while not requiring precise alignment information between text and speech. MaskGCT achieves high-quality text-to-speech synthesis using text to predict semantic tokens extracted from a speech self-supervised learning (SSL) model, and then predicting acoustic tokens conditioned on these semantic tokens. Our experiments demonstrate that MaskGCT outperforms the state-of-the-art TTS system on speech quality, similarity, and intelligibility with scaled model size and training data, and MaskGCT can control the total duration of generated speech. We also explore the scalability of MaskGCT in tasks such as speech translation, voice conversion, emotion control, and speech content editing, demonstrating the potential of MaskGCT as a foundational model for speech generation.
Appendix A Details of MaskGCT
Model Architecture. We use a Llama-style [Touvron et al., 2023] Transformer architecture as the backbone of our model, incorporating gated linear units with GELU [Hendrycks and Gimpel, 2016] activation (SwiGLU), rotation position encoding [Su et al., 2024], etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm [Zhang and Sennrich, 2019], which accepts the time step $t$ as the condition. Table 5 presents the key hyperparameters of the models.
Table 5: Overview of the key hyperparameters of MaskGCT.
T2S-Base T2S-Large S2A
Layers 16 16 16
Model Dimension 1,024 1,536 1,024
FFN Dimension 4,096 6,144 4,096
Attention Heads 16 16 16
Attention Type Bidirectional Bidirectional Bidirectional
Activation Function SwiGLU - -
Positional Embeddings RoPE ($\theta$ = 10,000) - -
Number of Parameters 315M 695M 353M
Inference Steps for the S2A model. The S2A model generates tokens layer by layer during inference. Since the acoustic codec follows an RVQ structure, we can view the S2A inference as a process from coarse to fine. We also use more iterations in the initial layers, as the first few layers carry more information. By default, we use inference steps of $[40,16,1,1,1,1,1,1,1,1,1,1]$ for each layer, however, we find that the S2A model can also perform well with fewer steps, such as $[10,1,1,1,1,1,1,1,1,1,1,1]$, with only a very slight performance loss.
Table 6: Evaluation results of different inference steps for the S2A model.
Inference Steps SIM-O $\uparrow$ WER $\downarrow$ FSD $\downarrow$
SeedTTS test-en
$[10,1,1,1,1,1,1,1,1,1,1,1]$ 0.709 2.796 0.164
$[40,16,1,1,1,1,1,1,1,1,1,1]$ 0.728 2.466 0.159
SeedTTS test-zh
$[10,1,1,1,1,1,1,1,1,1,1,1]$ 0.766 2.268 0.111
$[40,16,1,1,1,1,1,1,1,1,1,1]$ 0.777 2.183 0.101
Appendix B Details of Semantic/Acoustic Codec
For semantic codec, we train a VQVAE model using the hidden features from the 17th layer of W2v-BERT 2.0, incorporating factorized codec [Lezama et al., 2022] technology. The original hidden dimension of 1,024 is projected into a lower-dimensional space for quantization. The codebook size is set to 8,192, with a codebook dimension of 8. We employ only the $\mathcal{L}_{1}$ loss as the reconstruction target, optimizing the codebook with codebook loss and commitment loss. The input features are normalized to have a mean of 0 and a variance of 1, based on the statistics of the training dataset. The encoder and the decoder are each composed of 12 mirrored ConvNext blocks, featuring a kernel size of 7 and a hidden size of 384.
Figure 5: An overview of the semantic codec (left) and acoustic codec (right). The semantic codec is trained to quantize semantic features with a single codebook and reconstruct semantic features. The acoustic codec is trained to quantize and reconstruct the speech waveform using RVQ, with time and spectral discriminators to enhance the reconstruction quality further.
For acoustic codec, the basic architecture of the encoder follows Kumar et al. [2024] and the decoder follows Siuzdak [2023]. The Vocos-based decoder can model amplitude and phase, enabling waveform generation through inverse STFT transformation without requiring upsampling. The number of RVQ layers, codebook size, and codebook dimension are set to 12, 8,192, and 8, respectively. We utilize the multi-scale mel-reconstruction loss [Kumar et al., 2024] $\mathcal{L}_{\text{rec}}$, for the adversarial loss $\mathcal{L}_{\text{adv}}$, we employ both the multi-period discriminator (MPD) and the multi-band multi-scale STFT discriminator, as proposed by Lee et al. [2022], Kumar et al. [2024]. Additionally, we incorporate the relative feature matching loss $\mathcal{L}_{\text{feat}}$. For codebook learning, we use the codebook loss $\mathcal{L}_{\text{codebook}}$ and the commitment loss $\mathcal{L}_{\text{commit}}$ from VQ-VAE. We set $\lambda_{\text{rec}}=10.0$, $\lambda_{\text{adv}}=2.0$, $\lambda_{\text{feat}}=2.0$, $\lambda_{\text{codebook}}=1.0$, $\lambda_{\text{commit}}=0.25$ as coefficients for balancing each loss terms. Figure 5 shows the overview of the semantic codec and acoustic codec, Table 7 presents the detailed model configurations of semantic codec and acoustic codec.
Table 7: The detailed model configurations of semantic codec and acoustic codec.
Semantic Codec Acoustic Codec
Input W2v-BERT 2.0 hidden Waveform
Sample Rate 16K 24K
Hopsize 320 480
Number of (R)VQ Blocks 1 12
Codebook size 8,192 1,024
Codebook Dimension 8 8
Decoder Hidden Dimension 384 512
Decoder Kernel Size 7 7
Number of Decoder Blocks 12 30
Number of Parameters 44M 170M
Appendix C Details of Duration Predictor
MaskGCT requires specifying the target speech duration during inference, so we train a flow matching [Lipman et al., 2022, Liu et al., 2022b] based duration predictor to obtain the total duration of the target audio by summing the phone-level duration. Note that we do not need to actually use the phone-level durations but only use them to make a reasonable estimate of the total duration, leaving other total duration predictor methods for future works to explore. The duration predictor has a similar Transformer architecture to MaskGCT, with 12 layers, 12 attention heads, and a hidden size of 768. We also adapt in-context learning and classifier-free guidance for the duration predictor. During training, we randomly select a prefix segment of the phoneme sequence and its corresponding duration as a prompt, which is not added with noise. At the same time, we use a probability of 0.15 to drop the prompt. We model the duration in the log domain using flow matching. We denote $x_{1}$ as a random variable of $\log(\text{duration}+1)$, $x_{0}$ as a randomly sampled Gaussian noise, then $v_{\theta}(x_{t},t)=x_{t}=(1-t)x_{0}+tx_{1}$, where the timestep $t\in[0,1]$. The loss function of the duration predictor is $\mathbb{E}_{t,x_{1}}(v_{\theta}(x_{t},t)-(x_{1}-x_{0}))^{2}$. In the inference stage, we use a midpoint ODE solver to generate the target from randomly sampled Gaussian noise with a total of 4 steps. We pretrain a duration aligner (between phoneme and W2v-BERT 2.0 semantic feature) based on monotonic alignment search (MAS) [Kim et al., 2020] to get the ground truth duration for each phoneme.
Appendix D Text Tokenizer
We consider two text tokenization methods: Grapheme-to-Phoneme (G2P) and Byte Pair Encoding (BPE). For G2P, we employ phonemize  [^3] https://github.com/bootphon/phonemizer for English and a combination of jieba  [^4] https://github.com/fxsjy/jieba and pypinyin  [^5] https://github.com/mozillazg/python-pinyin for Chinese. For BPE, we utilize the BPE method and vocabulary from Whisper  [^6] https://github.com/huggingface/transformers/blob/main/src/transformers/models/whisper/tokenization_whisper.py, with a vocabulary size exceeding 30,000.
Appendix E Classifier-Free Guidance
We adopt the classifier-free guidance [Ho and Salimans, 2022] technique for both the T2S model and the S2A model. We also introduce classifier-free guidance with rescaling, following Lin et al. [2024]. In the training stage, we randomly drop the prompt with a probability of 0.15 to model the probability distribution $p_{\theta}(\mathbf{X})$ without the prompt. During inference, we compute the output embedding $g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g_{\theta}(\mathbf{X}|\mathbf{X}^{p})+w_{\text{cfg}}\cdot(g_{\theta}(\mathbf{X}|\mathbf{X}^{p})-g_{\theta}(\mathbf{X}))$ of the last layer of the model, where $w_{\text{cfg}}$ is the classifier-free guidance scale, then we compute the rescale embedding $g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})=g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})\times\text{std}(g_{\theta}(\mathbf{X}|\mathbf{X}^{p}))/\text{std}(g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p}))$, the final output embedding is computed as $w_{\text{rescale}}\cdot g^{\text{rescale}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})+(1-w_{\text{rescale}})\cdot g^{\text{cfg}}_{\theta}(\mathbf{X}|\mathbf{X}^{p})$. In our paper, $w_{\text{cfg}}$ and $w_{\text{rescale}}$ are set as 2.5 and 0.75 by default.
Appendix F Evaluation Baselines
VALL-E [Wang et al., 2023]. A large-scale TTS system uses an autoregressive and an additional non-autoregressive model to predict discrete tokens from a neural speech codec [Défossez et al., 2022]. We reproduce VALL-E with Amphion toolkit [Zhang et al., 2023] and Librilight [Kahn et al., 2020] dataset.
NaturalSpeech 3 [Ju et al., 2024]. A non-autoregressive model large-scale TTS systems with factorized speech codec for speech decoupling representation and factorized diffusion Models for speech generation. It achieves human-level naturalness on the LibriSpeech test set. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
VoiceBox [Le et al., 2024]. A non-autoregressive model large-scale multi-task speech generation model based on flow matching Lipman et al. [2022]. We report the scores of LibriSpeech test-clean obtained from Ju et al. [2024] and ask for the generated samples for subjective evaluation.
XTTS-v2 [Casanova et al., 2024]. An open-source multilingual TTS model that supports 16 languages. It is also based on an autoregressive model. We use the official code and pre-trained checkpoint  [^7] https://huggingface.co/coqui/XTTS-v2.
VoiceCraft [Peng et al., 2024]. A token-infilling neural codec language model for text editing and text-to-speech. It predicts multi-layer tokens in a delay pattern. We use the official code and pre-trained checkpoint  [^8] https://huggingface.co/pyp1/VoiceCraft/blob/main/830M_TTSEnhanced.pth.
CosyVoice [Du et al., 2024a]. A two-stage large-scale TTS system. The first stage is an autoregressive model and the second stage is a diffusion model. It is trained on 170,000 hours of multilingual speech data. We use the official code and pre-trained checkpoint  [^9] https://huggingface.co/model-scope/CosyVoice-300M.
Appendix G Post-Training for Emotion Control
MaskGCT can unlock more extensive capabilities with post-training. We take emotion control as an example. After being pretrained on a large-scale dataset, we fine-tune the T2S model by adding an additional emotion label as a prefix to the original input sequence. We use a small emotion dataset, ESD [Zhou et al., 2021], which consists of 350 parallel utterances with an average duration of 2.9 seconds spoken by 10 native English and 10 native Mandarin speakers, to fine-tune our model. The experimental results show that MaskGCT can unlock emotion control capabilities for zero-shot in-context learning scenarios. For the construction of the train and test datasets, we selected one male and one female speaker each from native English and native Mandarin backgrounds, resulting in a total of four speakers for the test dataset. The remaining 16 speakers were allocated to the train dataset. For the 350 parallel Chinese utterances, we randomly chose 22 utterances for the test set, with the remaining utterances designated for training. Similarly, for the 350 parallel English utterances, we randomly selected 21 utterances for the test set, with the rest used for training. To assess the consistency between the generated audio and the target emotion label, we trained an emotion classification model using the constructed train dataset. This model achieved a classification accuracy of 72% on the test dataset. We show some examples in our demo page.
Appendix H Speech Editing
Based on the mask-and-predict mechanism, our text-to-semantic model supports zero-shot speech content editing with the assistance of a text-speech aligner. By using the aligner, we can identify the editing boundary of the original semantic token sequence, mask the portion that needs to be edited, and then predict the masked semantic tokens using the edited text and the unmasked semantic tokens. However, we have observed that our system is not very robust in editing tasks. A possible conjecture is that we need to adopt a training paradigm better suited for editing tasks, such as fill-in-mask [Le et al., 2024, Du et al., 2024b]. We show some examples in our demo page.
Appendix I Voice Conversion
MaskGCT supports zero-shot voice conversion by fine-tuning the S2A with a modified training strategy. The zero-shot voice conversion task aims to alter the source speech to sound like that of a target speaker using a reference speech from the target speaker, without changing the semantic content. We can directly use the semantic tokens $\mathbf{S}_{\text{src}}$ extracted from the source speech and the prompt acoustic tokens $\mathbf{A}_{\text{ref}}$ extracted from the reference speech to predict the target acoustic tokens $\mathbf{A}_{\text{tgt}}$. Since $\mathbf{S}_{\text{src}}$ may retain some timbre information, we perform timbral perturbation on the semantic features input to the semantic codec encoder. Specifically, we apply timbral perturbation to the input mel-spectrogram features of the W2v-BERT 2.0 model, following the method outlined in FreeVC [Li et al., 2023b]. We fine-tune our S2A model using this training strategy. We show some examples in our demo page.
Appendix J Duration Controllable Speech Translation
The goal of the speech translation task is to translate speech from one language to another while preserving the original semantic, timbre, and prosody. In some scenarios, we also need to ensure that the total duration remains relatively unchanged, such as in video translation. Our model can achieve this seamlessly, with the ability to control the total duration and, through in-context learning, use the pre-translation speech as a prompt to maintain the timbre and prosody. We show some examples of speech translation between Chinese and English in our demo page.
Appendix K Hard Cases Evaluation
We evaluate the performance of MaskGCT on some hard cases (SeedTTS test-hard), which refer to instances where large-scale TTS models, particularly those AR-based models, often exhibit hallucinations. These cases include phrases with repeating words, tongue twisters, and other complex linguistic structures. Examples of such cases include: “the great greek grape growers grow great greek grapes”, “How many cookies could a good cook cook If a good cook could cook cookies? A good cook could cook as much cookies as a good cook who could cook cookies”, and “ thought a thought. But the thought I thought wasn’t the thought I thought I thought. If the thought I thought I thought had been the thought I thought, I wouldn’t have thought so much”.
Table 8: The evaluation results of MaskGCT and AR + SoundStorm on SeedTTS test-hard.
System SIM-O $\uparrow$ WER $\downarrow$
SeedTTS test-hard
AR + SoundStorm 0.692 34.16
AR + SoundStorm (rank 5) 0.739 17.05
MaskGCT 0.748 10.27
MaskGCT (rank 5) 0.776 6.258
Appendix L Discussion about Concurrent Works
SimpleSpeech [Yang et al., 2024], DiTTo-TTS [Lee et al., 2024], and E2 TTS [Eskimez et al., 2024] are also NAR-based models that do not necessitate precise alignment information between text and speech, nor do they forecast phoneme-level duration. These are concurrent works with MaskGCT. The three models all employ diffusion modeling on speech representations within continuous spaces. SimpleSpeech models the latent representation of a wav codec based on finite scalar quantization (FSQ) [Mentzer et al., 2023], DiTTo-TTS utilizes the latent representation of a wav codec based on residual vector quantization (RVQ), and E2 TTS directly models the mel-spectrogram with flow matching.
Appendix M Boarder Impact
Given that our model can synthesize speech with high speaker similarity, it carries potential risks of misuse, including spoofing voice identification or impersonating specific speakers. Our experiments were conducted under the assumption that the user consents to be the target speaker for speech synthesis. To mitigate misuse, it is essential to develop a robust model for detecting synthesized speech and to establish a system for reporting suspected misuse.